OpenStudy (anonymous):

MEDAL AND FAN : A right triangle has a base that measure 39 inches and a height that measures 80 inches. What is the length of its hypotenuse?

OpenStudy (kidrah69):

We need to use the Pythagorean theorem \[a^2+b^2=c^2\]

OpenStudy (anonymous):

Correct

OpenStudy (kidrah69):

right so do you know how to plug in the numbers?

OpenStudy (anonymous):

A^2+39^2=80^2 ?

OpenStudy (kidrah69):

its asking for the length of the hypotenuse you shouldnt have anything on c^2

OpenStudy (anonymous):

oh sorry so 39^2+80^2=_________

OpenStudy (kidrah69):

yep yep :D now solve that

OpenStudy (anonymous):

aka 1521+6400= ___________

OpenStudy (kidrah69):

wohoo well done good! :) now solve that

OpenStudy (anonymous):

that = 7921 and that is 89^2?

OpenStudy (kidrah69):

yep ::) good job! :D also welcome to openstudy!

OpenStudy (anonymous):

Thank you so much for both things :)))

OpenStudy (kidrah69):

haha you're welcome tag me if you need help @kidrah69 :)

OpenStudy (anonymous):

okay- i have more to solve and if you want to help you can :) I just need help setting up most of the time

OpenStudy (kidrah69):

okie report

OpenStudy (kidrah69):

REPOST :P